OpenStudy (anonymous):

Can you write 1 1/6 as a proper fraction? Explain.

OpenStudy (anonymous):

A mixed number is really an addition: 1 1/6 = 1 + 1/6 Write the 1 with a denominator of 6 and add: 6/6 +1/6 = 7/6 The quick way is to multiply the denominator times the whole number and add that product o the numerator: 1*6+1 = 7, then write that over the denominator of 6: 7/6

OpenStudy (anonymous):

\[Me + Ur Answer = Happiness :)\]

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

Can you re wright this in a simpler form? cuz i don't get wht u sayin!

OpenStudy (anonymous):

So is the answer yes.... 7/6?

OpenStudy (anonymous):

Is the answer yes or no?

OpenStudy (anonymous):

Hurry Please

OpenStudy (anonymous):

when you have a mixed number 1 1/6 it is read "one and one sixth." That means that a mixed number is really an addition of the whole number and the proper fraction:\[1 +\frac{1}{6}\]Tell me if its ok to continue. . .

OpenStudy (anonymous):

The answer is 7/6, yes

OpenStudy (anonymous):

Did you want further explanation?

OpenStudy (anonymous):

yes please cuz is tht a proper fraaction?

OpenStudy (anonymous):

I'll move on; if you want me to explain, respond and I'll come back :)

OpenStudy (anonymous):

and plz hurry

OpenStudy (anonymous):

Hello?

OpenStudy (anonymous):

Hello????????

OpenStudy (anonymous):

7/6 is not a proper fraction, and 1 1/6 can not be written as a proper fraction. I apologize that I did not read the question as closely as I should have.

OpenStudy (chriss):

A proper fraction is a part of 1 whole something. 7/6 as mentioned is not a proper fraction because it is 1whole something and 1 part out of 6 total parts of a second whole something. So the final answer is correct that 1 1/6 cannot be written as a proper fraction as it is greater than one.